# SCAM: A Real-World Typographic Robustness Evaluation for Multimodal Foundation Models

### v4

@@ Abstract
Typographic attacks exploit the interplay between text and visual content in multimodal foundation models, causing misclassifications when misleading text is embedded within images. However, existing datasets are limited in size and diversity, making it difficult to study such vulnerabilities. In this paper, we introduce SCAM, the largest and most diverse dataset of real-world typographic attack images to date, containing 1,162 images across hundreds of object categories and attack words. Through extensive benchmarking of Vision-Language Models (VLMs) on SCAM, we demonstrate that typographic attacks significantly degrade performance, and identify that training data and model architecture influence the susceptibility to these attacks. Our findings reveal that typographic attacks persist in state-of-the-art Large Vision-Language Models (LVLMs) due to the choice of their vision encoder, though larger Large Language Models (LLMs) backbones help mitigate their vulnerability. Additionally, we demonstrate that synthetic attacks closely resemble real-world (handwritten) attacks, validating their use in research. Our work provides a comprehensive resource and empirical insights to facilitate future research toward robust and trustworthy multimodal AI systems. We publicly release the datasets introduced in this paper along with the code for evaluations at www.bliss.berlin/research/scam.
@@
1 Introduction
Vision-Language Models (VLMs) such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9]. With the growing use of VLMs in real-world applications, such as medical applications [10, 11, 12] and autonomous systems [13], as well as in Large Vision-Language Models (LVLMs) [14, 15, 16], it is essential to ensure the safety and reliability of these models. However, recent research has revealed that these models are vulnerable to typographic attacks [17, 18, 19], which exploit the reliance on the textual component of multimodal learning.
Typographic attacks target multimodal foundation models by inserting human-readable text with misleading semantics into an image, leading to incorrect model predictions. Figure 1 depicts how adding the word “taxi” to an image of a “clock” can cause a model to misclassify it as a “taxi”. These attacks expose a fundamental interaction in multimodal learning: the model gets tricked by focusing on text within an image, more than the visual content itself [17, 18, 19]. Understanding and mitigating such vulnerabilities is essential for the safe deployment of multimodal foundation models.
To facilitate research in this field, we introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest real-world typographic attack dataset to date. Each image contains an object with a semantically unrelated handwritten attack word [^1] While we use the term “word”, our attacks also incorporate phrases like “no left turn”. on a post-it note. Furthermore, we provide two variations, a clean dataset where attack words are removed (NoSCAM), and a synthetic dataset where attack words are digitally reintroduced (SynthSCAM).
We provide extensive evaluations over a range of VLMs and LVLMs and show that:
@@ 1 Introduction
    Models of both classes experience a significant drop in accuracy when typographic attacks are introduced (SCAM).
@@
    We find that synthetic attacks closely align with real-world attacks, providing empirical support for the established practice of evaluating on synthetic datasets.
@@
    LVLMs, such as the LLaVA family, inherit typographic vulnerabilities from their vision encoders.
@@
    Utilizing larger Large Language Model (LLM) backbones can alleviate an LVLM’s inherited weaknesses regarding typographic robustness.
Table 1: Performance of VLMs and LVLMs available through OpenCLIP resp. ollama and OpenAI on the SCAM datasets. Note that internally all LLaVA models that we evaluate utilize ViT-L-14-336 trained by OpenAI for the image encoding. Furthermore, ViT-bigG-14 trained on laion2b is used in the Kandinsky diffusion model [20]. This table presents a selection of VLMs; the full list of all 99 can be found in LABEL:tab:full-overview in the appendix.
@@
NoSCAM SCAM
RN50 openai 97.76 36.61 ↓61.15
ViT-B-32 laion2b 98.45 74.68 ↓23.77
ViT-B-16 laion2b 98.71 69.16 ↓29.55
ViT-B-16-SigLIP webli 99.22 81.40 ↓17.82
ViT-L-14 commonpool_xl 99.48 74.68 ↓24.80
openai 99.14 40.14 ↓59.00
ViT-L-14-336 openai 99.22 33.85 ↓66.37
ViT-g-14 laion2b 99.05 61.93 ↓37.12
ViT-bigG-14 laion2b 99.40 70.89 ↓28.51
llava-llama3-1.1:8b - 98.09 39.50 ↓58.59
llava-1.6:7b - 97.50 58.43 ↓39.07
llava-1.6:13b - 98.88 58.00 ↓40.88
llava-1.6:34b - 98.97 84.85 ↓14.11
llama3.2-vision:90b - 98.88 71.01 ↓27.87
gpt-4o-mini-2024-07-18 - 99.40 84.68 ↓14.72
gpt-4o-2024-08-06 - 99.48 96.82 02.67
@@ 2 Related Work
Multimodal pretraining, as established by CLIP [1] and SigLIP [2], leverages image-text pairs to capture rich semantics suitable for a variety of applications [21, 22]. This approach commonly features a contrastive objective to align visual and textual content in a shared latent space. Following this paradigm, various Vision-Language Models (VLMs) [23, 24, 25] have been proposed, demonstrating impressive performance in numerous downstream tasks [5, 6, 7]. Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks [26, 27, 28, 29, 28]. Due to their strong capabilities and versatility, VLM s and LVLM s function as the backbone for a wide range of domains such as text-to-image generation [8, 9].
@@
To study the intricacies of typographic attacks and evaluate the effectiveness of defense mechanisms, the majority of prior works [33, 34, 18, 19, 30, 35] synthetically add attack words into existing image datasets while only few [36, 37, 38] propose datasets that exhibit real-world typographic attacks. Existing real-world datasets, such as PAINT [38] and Materzynska+[37], lack the diversity and scale needed to reliably evaluate the vulnerabilities of multimodal foundation models. The previously largest real-world typographic attack dataset, RTA-100 [36], consists of 1,000 images spanning 100 objects and attack words. However, existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see table 2 for a comparison). Furthermore, current datasets lack a structured comparison between real-world and synthetic typographic attacks, preventing a systematic analysis of transferability between these domains. This gap hinders not only the rigorous evaluation of VLM s and LVLM s but also the development of effective defense mechanisms.
Figure 2: Examples of images from the SCAM dataset. Additional images for both NoSCAM and SynthSCAM are provided in fig. 7 in the appendix.
@@
3 SCAM Dataset
We introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest and most diverse real-world typographic attack dataset to date, containing 1,162 images (see table 2). The dataset and its two variations (NoSCAM and SynthSCAM), allowing for an empirical assessment of transferability between these settings, are publicly available on Hugging Face, along with the fully compatible and publicly accessible SCAM benchmark repository for standardized evaluation.
Data Collection.
We collected 1,162 images from diverse environments, such as households, utility stores, and traffic infrastructure, ensuring a varied and representative dataset. An overview of example images from SCAM is presented in fig. 2. Each image in the SCAM dataset captures an object together with an attack word – unrelated in meaning – handwritten on a single post-it note in English and placed alongside the object, as illustrated in fig. 1a.
To create the cleaned NoSCAM dataset, we remove all attack words by overlaying the post-it notes with their original color, thereby obscuring the text while preserving the original object images. Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font), matching the color of the ink from the original images.
Data Annotation and Processing.
[…]
Dataset Statistics.
SCAM comprises 1,162 images, featuring 660 distinct object labels and 206 unique attack words, that form 1,147 unique object-word combinations. We provide lists of the 660 unique object labels and 206 attack words in section A.2. We categorize the attack words into ten groups, covering both everyday words and safety-critical terms – such as “pedestrian” or “no left turn” – that are essential for autonomous driving systems. The distribution of these groups is depicted in fig. 3.
As shown in table 2, SCAM significantly surpasses existing real-world datasets in diversity, both in terms of object counts and attack words. Compared to RTA-100 [36], SCAM covers a much broader range of objects (660 vs. 100) and attack words (206 vs. 100), offering a more comprehensive evaluation of model robustness to typographic attacks. This increased diversity makes SCAM a stronger benchmark for assessing the vulnerabilities of multimodal foundation models.
@@ 3 SCAM Dataset
Dataset Images Objects Attack words
@@
RTA-100 [36] 1,000 100 100
SCAM (ours) 1,162 660 206
@@ 4 Evaluation
To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [41, 42, 1, 43, 44]. We also evaluate seven LVLMs, including GPT-4o [26], LLaVA [27, 29, 28] and Llama3.2-vision [45] models. The LLaVA and Llama models are deployed via the Ollama platform [46], whereas GPT-4o is accessed via OpenAI’s API. An overview of selected models is provided in table 1, while LABEL:tab:full-overview presents the complete list of all 106 evaluated models.
@@
We evaluate the performance of VLMs in a zero-shot classification task on the SCAM datasets. For each image, we compute the cosine similarity between its embedding and the text embeddings of both the object label and the attack word. These text embeddings are generated using the set of text templates proposed in [1], as detailed in section A.3. The predicted label is then determined based on the highest cosine similarity score. An overview of this process is depicted in fig. 1b. Here, when the model is presented with an image of a “clock” with the attack word “taxi”, it does not classify the image correctly. All evaluations of VLMs were performed on an A100 40GB card.
@@
To assess the robustness of an LVLM against typographic attacks, we evaluate whether its output changes when exposed to typographic modifications. Our evaluation follows the setup illustrated in fig. 1c, inspired by [47]. The model is provided with an image from SCAM, NoSCAM, or SynthSCAM, along with the following prompt:
@@
Respond strictly with only (a) or (b), nothing else.
@@ 5 Results
We demonstrate that multimodal foundation models are susceptible to typographic attacks in real-world data, as shown in table 1. Our results show that typographic attacks on VLMs lead to an average drop in accuracy of 26 percentage points. LVLMs exhibit similar behavior and we find that the choice of vision encoder plays a critical role in determining the extent of accuracy degradation.
Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig. 4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table 3. These results show that synthetic typographic attacks closely mirror real-world scenarios, making them a reliable proxy for evaluating the robustness of multimodal foundation models.
@@
Table 3: Alignment between real-world (SCAM) and synthetic (SynthSCAM) attacks. The tables present average confusion matrices. The similarity in outcomes indicates that synthetic attacks closely replicate the effects of real attacks. Values represent the mean ± standard deviation, rounded to integers.
((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
[…]
((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
[…]
Zero-Shot Results for VLMs.
As shown in fig. 4, VLMs achieve high prediction accuracy in the zero-shot classification task on the NoSCAM dataset. However, accuracy declines for almost all models when evaluated on the SCAM and SynthSCAM datasets. Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 36 percentage points. For a detailed breakdown of model accuracies across the three datasets, see LABEL:tab:full-overview in the appendix.
Impact of the Training Dataset and Model Architecture
Table 4 presents the impact of training data on the ViT-B models [48] trained by the OpenCLIP team. We focus solely on these models as they share the same training paradigm but are trained on different datasets. Overall, the ViT-B models trained on the LAION datasets [41] are highly susceptible to typographic attacks, while certain CommonPool variants [49] exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively. Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 [1] model from OpenAI. Consequently, as indicated by the accuracies for that model in table 1, this filtering criterion appears to propagate susceptibility.
@@
Table 5: Comparison of average accuracy on SCAM for different CLIP and SigLIP models trained by OpenAI and Visheratin [50] on LAION-COCO NLLB (referred to, in OpenCLIP, as v1). More details about the selection process can be found in section A.5.
Training data Architecture Accuracy %
@@
While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance. We compare different architectures: CLIP-RN (ResNets [51]) and CLIP-ViT models trained by OpenAI [1], as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB [50] (see table 5). Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models. Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
Our evaluations show that the robustness of VLMs against typographic attacks varies with patch size, as illustrated in table 6. Notably, smaller patch sizes generally result in higher susceptibility to typographic attacks.
Table 6: Models become more susceptible to typographic attacks as patch size decreases. Dataset fixed to laion2b, models evaluated on SCAM.
Model Patch Size Accuarcy %
@@
Lastly, fig. 6 shows that there is no noticeable correlation between model size and accuracy drop caused by typographic attacks. However, we find that the original CLIP models [1] trained by OpenAI perform significantly worse across all sizes.
Figure 5: Susceptibility to typographic attack (evaluation on SCAM) is agnostic of VLMs size, measured in millions of parameters.
Figure 6: Model accuracy on SCAM decreases as the post-it area increases. Shaded area shows mean ± standard deviation across all 99 VLMs that we evaluate.
@@
We observe a strong correlation between zero-shot accuracy and post-it area, indicating that the size of the attack text affects accuracy (see fig. 6). The mean accuracy across all models follows this trend, which is consistent with the findings in [34].

### v6

@@ Abstract
Typographic attacks exploit the interplay between text and visual content in multimodal foundation models, causing misclassifications when misleading text is embedded within images. Existing datasets are limited in size and diversity, making it difficult to study such vulnerabilities. In this paper, we introduce SCAM, the largest and most diverse dataset of real-world typographic attack images to date, containing 1162 images across hundreds of object categories and attack words. Through extensive benchmarking of Vision-Language Models on SCAM, we demonstrate that typographic attacks significantly degrade performance, and identify that training data and model architecture influence the susceptibility to these attacks. Our findings indicate that typographic attacks remain effective against state-of-the-art Large Vision-Language Models, especially those employing vision encoders inherently vulnerable to such attacks. However, employing larger Large Language Model backbones reduces this vulnerability while simultaneously enhancing typographic understanding. Additionally, we demonstrate that synthetic attacks closely resemble real-world (handwritten) attacks, validating their use in research. Our work provides a comprehensive resource and empirical insights to facilitate future research toward robust and trustworthy multimodal AI systems. Finally, we publicly release the datasets introduced in this paper, along with the code for evaluations under www.bliss.berlin/research/scam.
1 Introduction
Vision-Language Models (VLMs) such as CLIP [1] and SigLIP [2] have demonstrated remarkable performance across various multimodal tasks [3], enabling applications in image classification [1, 4], multimodal retrieval [5, 6], and generative tasks [7, 8, 9]. However, recent research has revealed that these models are vulnerable to typographic attacks [10, 11, 12], which exploit the reliance on the textual component of multimodal learning.
Typographic attacks target multimodal foundation models by inserting human-readable text with misleading semantics into an image, leading to incorrect model predictions. Figure˜1 depicts how adding the word “taxi” to an image of a “clock” can cause a model to misclassify it as a “taxi”. These attacks expose a fundamental interaction in multimodal learning: the model gets tricked by focusing on text within an image, more than the visual content itself. Understanding and mitigating such vulnerabilities is essential for the safe deployment of multimodal foundation models.
@@ 1 Introduction
With the growing use of VLMs in real-world applications, such as medical applications [13, 14, 15] and autonomous systems [16], as well as in Large Vision-Language Models (LVLMs) [17, 18, 19], it is essential to ensure the safety and reliability of these models.
To facilitate research in this field, we introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest and most diverse real-world typographic attack dataset to date. Each image contains an object with a semantically unrelated handwritten attack word [^1] While we use the term “word”, our attacks also incorporate phrases like “no left turn”. on a post-it note. Furthermore, we provide two variations, a clean dataset where attack words are removed (NoSCAM), and a synthetic dataset where attack words are digitally reintroduced (SynthSCAM).
In summary, we provide extensive evaluations over a range of VLMs and LVLMs and show that:
@@
    Models of both classes experience a significant drop in accuracy when faced with real-world typographic attacks (SCAM).
@@
    Synthetic attacks (SynthSCAM) closely align with real-world attacks, providing empirical support for the established practice of evaluating on synthetic datasets.
@@
    LVLMs, such as the LLaVA family, exhibit vulnerabilities to typographic attacks, reflecting weaknesses inherited from their vision encoders.
@@
    Employing larger Large Language Model (LLM) backbones reduces susceptibility to typographic attacks in LVLMs while also improving their typographic understanding.
Table 1: Performance of VLMs and LVLMs available through OpenCLIP [20] resp. ollama [21], and Anthropic’s resp. OpenAI’s API on the SCAM datasets. Note that internally all LLaVA models that we evaluate utilize ViT-L-14-336 trained by OpenAI for the image encoding. Furthermore, ViT-bigG-14 trained on laion2b is used in the Kandinsky diffusion model [22]. This table presents a only selection of VLMs; the full list of all 99 can be found in LABEL:tab:full-overview in the appendix.
@@
NoSCAM SCAM SynthSCAM
RN50 openai 97.76 36.61 $\downarrow$61.15 22.39 $\downarrow$75.37
ViT-B-32 laion2b 98.45 74.68 $\downarrow$23.77 56.16 $\downarrow$42.29
ViT-B-16 laion2b 98.71 69.16 $\downarrow$29.55 52.28 $\downarrow$46.43
ViT-B-16-SigLIP webli 99.22 81.40 $\downarrow$17.82 75.45 $\downarrow$23.77
ViT-L-14 commonpool_xl 99.48 74.68 $\downarrow$24.80 68.56 $\downarrow$30.92
openai 99.14 40.14 $\downarrow$59.00 26.61 $\downarrow$72.53
ViT-L-14-336 openai 99.22 33.85 $\downarrow$65.37 26.44 $\downarrow$72.78
ViT-L-14-CLIPA-336 datacomp1b 99.57 74.76 $\downarrow$24.81 65.81 $\downarrow$33.76
ViT-g-14 laion2b 99.05 61.93 $\downarrow$37.12 47.03 $\downarrow$52.02
ViT-bigG-14 laion2b 99.40 70.89 $\downarrow$28.51 58.91 $\downarrow$40.49
llava-llama3:8b - 98.09 39.50 $\downarrow$58.59 44.04 $\downarrow$54.05
llava:7b-v1.6 - 97.50 58.43 $\downarrow$39.07 61.45 $\downarrow$36.06
llava:13b-v1.6 - 98.88 58.00 $\downarrow$40.88 55.42 $\downarrow$43.46
llava:34b-v1.6 - 98.97 84.85 $\downarrow$14.11 85.11 $\downarrow$13.86
gemma3:4b - 97.24 58.05 $\downarrow$39.19 64.80 $\downarrow$32.44
gemma3:12b - 99.14 52.02 $\downarrow$47.12 64.80 $\downarrow$34.34
gemma3:27b - 97.42 81.67 $\downarrow$15.75 83.65 $\downarrow$13.77
llama3.2-vision:90b - 98.88 71.01 $\downarrow$27.87 74.65 $\downarrow$24.22
llama4:scout - 99.23 88.12 $\downarrow$11.10 87.18 $\downarrow$12.05
gpt-4o-mini-2024-07-18 - 99.40 84.68 $\downarrow$14.72 87.09 $\downarrow$12.31
claude-sonnet-4-20250514 - 99.31 91.13 $\downarrow$8.18 90.28 $\downarrow$9.03
gpt-4o-2024-08-06 - 99.48 96.82 $\downarrow$2.67 94.58 $\downarrow$4.90
@@ 2 Related Work
Multimodal pretraining, as established by CLIP [1] and SigLIP [2], leverages image-text pairs to capture rich semantics suitable for a variety of applications [23, 24]. This approach commonly features a contrastive objective to align visual and textual content in a shared latent space. Following this paradigm, various Vision-Language Models (VLMs) [25, 26, 27] have been proposed, demonstrating impressive performance in numerous downstream tasks [5, 6, 7]. Recent research on Large Vision-Language Models (LVLMs) aligns the strong abilities of Large Language Models (LLMs) and vision encoders of VLMs to attain strong performance in vision-language tasks [28, 29, 30, 31, 32, 33, 34, 35].
Due to their strong capabilities and versatility, VLM s and LVLM s function as the backbone for a wide range of domains such as text-to-image generation [8, 9].
@@
To study the intricacies of typographic attacks and evaluate the effectiveness of defense mechanisms, the majority of prior works [39, 40, 11, 12, 36, 41] synthetically add attack words into existing image datasets while only few [42, 43, 44] propose datasets that exhibit real-world typographic attacks. Existing real-world datasets, such as PAINT [42] and Materzynska+[43], lack the diversity and scale needed to reliably evaluate the vulnerabilities of multimodal foundation models. The previously largest real-world typographic attack dataset, RTA-100 [44], consists of 1000 images spanning 100 objects and attack words. However, these existing real-world datasets remain limited in both scale and diversity, restricting their ability to comprehensively assess the vulnerabilities of multimodal foundation models (see LABEL:tab:dataset_comparison for a comparison). Our dataset addresses this limitation by introducing substantially larger variability across object–word combinations, which mitigates dataset bias and enables a more reliable assessment of generalization in VLMs and LVLMs. Furthermore, current datasets lack a structured comparison between real-world and synthetic typographic attacks, preventing a systematic analysis of transferability between these domains. SCAM is the first dataset to provide three explicitly aligned components: (i) SCAM, containing real-world attacked scenes, (ii) NoSCAM, the corresponding clean counterparts of the exact same scenes, and (iii) SynthSCAM, synthetically manipulated versions of the same object–word pairs. This tri-part design enables precise, controlled measurement of typographic attack effects and systematic comparisons across clean, real, and synthetic domains, establishing a stronger foundation for rigorous evaluation of VLMs and LVLMs as well as the development of effective defense mechanisms.
Figure 2: Cropped examples from the SCAM dataset, illustrating the natural variability introduced by nine different contributors using distinct smartphones, handwriting styles, and recording environments. This diversity results in variations in lighting, camera quality, backgrounds, and writing appearance. Additional uncropped examples for both NoSCAM and SynthSCAM are provided in fig.˜7 in the appendix.
3 SCAM Dataset
We introduce Subtle Character Attacks on Multimodal Models (SCAM), the largest and most diverse real-world typographic attack dataset to date, containing 1162 images (see LABEL:tab:dataset_comparison). The dataset and its two variations (NoSCAM and SynthSCAM), enabling empirical assessment of transferability between these settings, are publicly available on Hugging Face together with the fully compatible and publicly accessible SCAM benchmark repository for standardized evaluation.
Data Collection.
SCAM was collaboratively constructed by nine contributors using nice diffrent smartphones, ensuring natural variation in image quality. The data were recorded across diverse indoor and outdoor environments, including households, utility stores, and traffic infrastructure, under heterogeneous lighting conditions and viewing angles. This variability, illustrated in fig.˜2, reflects real-world conditions more faithfully than prior datasets.
Each image in the SCAM dataset captures a physical object together with an unrelated attack word handwritten in English on a single post-it note and placed next to the object, as shown in fig.˜1a. Sticky notes were deliberately chosen as the attack medium because they are common in everyday contexts (e.g., labeling in offices, classrooms, and shops) and provide a realistic, reproducible surface for typographic attacks.
To create the cleaned NoSCAM dataset, we remove all attack words by overlaying the post-it notes with their original color, thereby obscuring the text while preserving the original object images. This provides clean, paired counterparts for every attacked image. Additionally, we create the synthetic SynthSCAM dataset by reintroducing the original attack words into the images (using the Roboto font), matching the color of the ink from the original images. Choosing a uniform font style in this synthetic variant was a deliberate simplification. Prior synthetic attack datasets have already explored variations in font position, opacity, and color [40], whereas our focus is on providing clean, attacked, and synthetic counterparts of the same scenes.
Data Annotation and Processing.
All images were manually annotated with object and attack word labels as well as post-it size (in %) as part of a voluntary community effort. The annotation process was conducted using Label Studio [45]. To standardize the dataset, all images were zero-padded and downscaled to $512\times 512$ pixels using cv2.INTER_AREA [46], an interpolation method that preserves details while minimizing distortion.
Dataset Statistics.
SCAM comprises 1162 images, featuring 660 distinct object labels and 206 unique attack words, that form 1,147 unique object-word combinations. The complete lists of objects and attack words are provided in section˜A.8. We categorize the attack words into ten groups, covering both everyday words and safety-critical terms – such as “pedestrian” or “no left turn” – that are essential for autonomous driving systems. The distribution of these groups is depicted in fig.˜3.
@@ 3 SCAM Dataset
Dataset Images Objects Attacks
@@
RTA-100 [44] 1000 100 100
SCAM (ours) 1162 660 206
As summarized in LABEL:tab:dataset_comparison, SCAM significantly surpasses existing real-world datasets in diversity, both in terms of object counts and attack words. Compared to RTA-100 [44], SCAM covers a much broader range of objects (660 vs. 100) and attack words (206 vs. 100), offering a more comprehensive evaluation of model robustness to typographic attacks. The three explicitly aligned components – SCAM, NoSCAM, and SynthSCAM – allow precise, controlled measurement of typographic attack effects and direct comparisons between clean, real, and synthetic conditions, thereby closing critical gaps in prior datasets.
@@ 4 Evaluation
To evaluate VLMs such as CLIP [1] and SigLIP [2], we use the OpenCLIP suite [47, 48, 1, 49, 20]. We also evaluate eleven LVLMs, including Claude Sonnet 4 [35], GPT-4o [28], LLaVA [29, 30, 31], Gemma3 [32], Llama3.2-vision [33] and Llama4 [34] models. The LLaVA, Gemma3, and Llama models are deployed via the Ollama platform [21], whereas Claude Sonnet 4 and GPT-4o is accessed via OpenAI’s resp. Anthropic’s API. An overview of selected models is provided in table˜1, while LABEL:tab:full-overview presents the complete list of all 111 evaluated models.
@@
We evaluate the performance of VLMs in a zero-shot classification task. For each image, we compute the cosine similarity between its embedding and the text embeddings of both the object label and the attack word. These text embeddings are generated using the set of text templates proposed in [1], as detailed in section˜A.2. The predicted label is then determined based on the highest cosine similarity score. An overview of this process is depicted in fig.˜1b. Here, when the model is presented with an image of a “clock” with the attack word “taxi”, it does not classify the image correctly. All evaluations of VLMs were performed on an A100 40GB card.
@@
To assess the robustness of an LVLM against typographic attacks, we evaluate whether its output changes when exposed to typographic modifications. Our evaluation follows the setup illustrated in fig.˜1c, inspired by [50]. The model is provided with an image along with the following prompt:
@@
Respond strictly with only (a) or (b),
nothing else.
In this setup, the model correctly outputs “(a)” (the object label) when no attack word is present, as in the NoSCAM dataset. However, when the attack word is included, the model changes its response to “(b)”.
@@
All LLaVA and Gemma3 models were evaluated on an A100 40GB card, and llama3.2-vision:90b and llama4:scout on an H100 80GB.
@@ 5 Results
We demonstrate that multimodal foundation models are susceptible to typographic attacks in real-world data, as shown in table˜1. Our results show that typographic attacks on VLMs lead to an average drop in accuracy of 26 percentage points. LVLMs exhibit similar behavior, with the choice of vision encoder appearing to influence the extent of accuracy degradation. Furthermore, our analysis reveals that typographic attacks in real-world and synthetic data similarly affect model accuracy, as demonstrated by the accuracy distributions in fig.˜4 and the alignment between SCAM and SynthSCAM reflected in the confusion matrices in table˜4. These results show that synthetic typographic attacks closely mirror real-world scenarios, making them a reliable proxy for evaluating the robustness of multimodal foundation models.
@@
Table 4: Alignment between real-world (SCAM) and synthetic (SynthSCAM) attacks. The tables present average confusion matrices. The similarity in outcomes indicates that synthetic attacks closely replicate the effects of real attacks. Values represent the mean $\pm$ standard deviation, rounded to integers.
((a)) Average across all VLMs with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
SCAM fails $648\pm 192$ $154\pm 68$
SCAM succeeds $18\pm 11$ $341\pm 151$
((b)) Average across all LVLMs. with over 90% accuracy on NoSCAM.
SynthSCAM SynthSCAM
fails succeeds
SCAM fails $741\pm 243$ $96\pm 47$
SCAM succeeds $124\pm 71$ $201\pm 137$
Zero-Shot Results for VLMs.
As shown in fig.˜4, VLMs achieve high prediction accuracy in the zero-shot classification task on the NoSCAM dataset. However, accuracy declines for almost all models when evaluated on the SCAM and SynthSCAM datasets. Specifically, accuracy on SCAM drops by approximately 26 percentage points on average, while the SynthSCAM dataset exhibits an even steeper decline of 36 percentage points. For a detailed breakdown of model accuracies across the three datasets, see LABEL:tab:full-overview in the appendix.
Impact of Training Dataset and Model Architecture.
Table 5: VLMs’ architecture impacts susceptibility to typographic attacks. SigLIP outperforms CLIP, ViT outperforms ResNets, and ViTs become more susceptible to typographic attacks as patch size decreases. The tables present (average) accuracy (Acc.) % on SCAM.
((a)) Average across different CLIP and SigLIP models trained by OpenAI and Visheratin [51] on LAION-COCO NLLB (referred to, in OpenCLIP, as v1). More details about the selection process can be found in section˜A.4.
Training data Architecture Acc. %
@@
((b)) Dataset fixed to laion2b.
Model Patch Size Acc. %
@@
Table˜3 presents the impact of training data on the ViT-B models [52] trained by the OpenCLIP team. We focus solely on these models as they share the same training paradigm but are trained on different datasets. Overall, the ViT-B models trained on the LAION datasets [47] are highly susceptible to typographic attacks, while certain CommonPool variants [53] exhibit lower vulnerability, such as the text and image filtered datasets, generated based on textual and visual overlaps with ImageNet classes, respectively. Notably, the vulnerable commonpool_l_clip dataset results from selecting data points based on cosine similarity scores computed from image and text embeddings produced by the pretrained ViT-L-14 [1] model from OpenAI. Consequently, as indicated by the accuracies for that model in table˜1, this filtering criterion appears to propagate susceptibility.
While training data significantly affects robustness, it is equally important to understand how the model architecture impacts typographic attack resistance. We compare different architectures: CLIP-RN (ResNets [54]) and CLIP-ViT models trained by OpenAI [1], as well as CLIP-ViT and SigLIP-ViT models trained on LAION-COCO NLLB [51] (see table˜5(a)). Our findings show that, on average, CLIP-ViT models exhibit greater resilience to typographic attacks compared to CLIP-RN models. Similarly, SigLIP-ViT shows increased robustness compared to CLIP-ViT.
Moreover, our evaluations show that the robustness of VLMs against typographic attacks varies with patch size, as illustrated in table˜5(b). Notably, smaller patch sizes generally result in higher susceptibility to typographic attacks.
Lastly, fig.˜6 shows that there is no noticeable correlation between model size and accuracy drop caused by typographic attacks. However, we find that the original CLIP models [1] trained by OpenAI perform significantly worse across all sizes.
@@
We observe a strong correlation between zero-shot accuracy and post-it area, indicating that the size of the attack text affects accuracy (see fig.˜6). The mean accuracy across all models follows this trend, which is consistent with the findings in [40].
Figure 5: Susceptibility to typographic attack (accuracy on SCAM) is agnostic of VLM size, measured in millions of parameters.
Figure 6: Model accuracy on SCAM decreases as the post-it area increases. Shaded area shows mean $\pm$ standard deviation across all 99 VLMs that we evaluate.